# The Fragile Nature of Road Transportation Systems

### v2

@@ 4.2 Supply disruption
Because all the four components of the Eq. 66 have been demonstrated above to be non-negative, thus $d^{2}TTS/dr^{2}$ is also non-negative and we’ve proven the fragile nature of road transportation systems regarding the recovery process of supply disruptions.
@@ 5 Numerical simulation
Even though researchers generally consider MFDs to be well-defined, road transportation systems in the real world and their MFDs are always subject to stochasticity all the time, as shown in Geroliminis and Daganzo (2008); Saffari et al. (2022); Ambühl et al. (2021), and this is the case for FDs as well (Qu et al., 2017; Siqueira et al., 2016). Therefore, when validating a newly proposed traffic control algorithm, it has become a common practice to account for model uncertainties and showcase the method’s robustness, such as in Geroliminis et al. (2013); Haddad and Mirkin (2017); Zhou and Gayah (2023). In our study, however, the model stochasticity cannot be directly reflected in the mathematical analysis, hence, it is indispensable to show the influence of realistic stochasticity on the fragile nature of transportation systems with a numerical simulation, i.e., whether the system still maintains the same fragile response under real-world errors in MFD when a demand or supply disruption is present.
@@
In this section, we simulate the disruption recovery process. The MFD of the studied region is generated by applying MoC following Daganzo and Geroliminis (2008) with realistic parameters in the city center of Zurich. Some parameters, e.g., free-flow speed, back-propagation speed, maximal density, and capacity are provided in Ambühl et al. (2020) for Zurich with queried routes in Google API and with other validation methods. The total and average lane length for the city center is determined through SUMO with OpenStreetMap API. The average trip length of Zurich is studied in Schüssler and Axhausen (2008). We introduce stochasticity in the city center of Zurich with real traffic light data, which is publicized by the Statistical Office of Zurich and accessible in Genser et al. (2023). The authors acknowledge that MoC is developed with the premise of a homogeneous region, and given that the available data is limited to only one main intersection in this region, we assume that this intersection serves as a representative sample for the city center region. Also, since the signalization in Zurich is actuated based on the present traffic flow, they do not strictly follow a fixed-time signal cycle. Despite this actuation, a concentrated distribution can be easily observed in the dataset and we assume the green split of the cycle follows a normal distribution. The offset is considered to be zero in a way similar to the actuated signal in Yokohama in Daganzo and Geroliminis (2008). According to the daily average traffic density of Zurich in Ambühl et al. (2021), we approximate the traffic demand, which is also the trip completion when the traffic state is at equilibrium, is about $0.6veh/s$ for our studied region. This corresponds to an accumulation of around 975 vehicles in the city center. The parameters are summarized in Tab. 2.
Table 2: Estimated parameters for the city center of Zurich
Parameters
Notation
Unit
Value
Free-flow speed
$u_{0}$
$\rm m/s$
12.5
Back-propagation speed
$w_{0}$
$\rm m/s$
6.0
Maximal density
$\kappa$
$\rm veh/m$
0.145
Capacity
$s$
$\rm veh/s$
0.51
Total lane length
$D$
$\rm m$
68631
Average lane length
$l$
$\rm m$
167
Average trip length
$L$
$\rm m$
7110
Signal cycle time
$C$
$\rm s$
50
Signal green time (mean)
$\mu_{G}$
$\rm s$
14.8
Signal green time (std.)
$\sigma_{G}$
$\rm s$
2.5
Offset
$\delta$
$\rm s$
0
Traffic demand
$q$
$\rm veh/s$
0.6
As the average green time of the signal is 14.8 s and its standard deviation is 2.5 s, following the MoC described in Daganzo and Geroliminis (2008), we can produce three groups of cuts with the interval of one standard deviation, i.e., the green time being $\mu_{G}-\sigma_{G}$, $\mu_{G}$, or $\mu_{G}+\sigma_{G}$, and each group of cuts contains at least three cuts, as Fig. 9 shows. The group of cuts with a longer green time of signalization yields a greater MFD, and vice versa. The MFDs share the same maximal vehicle density of around 0.145 veh/m, which corresponds to a maximal accumulation of about 10000 vehicles for the studied region.
Figure 9: The MFD of the city center of Zurich through MoC

### v4

@@ 4.2 Supply disruption
Because all the four components of the Eq. 46 have been demonstrated above to be non-negative, thus $\frac{d^{2}TTS}{dn^{\prime 2}}$ is also non-negative and we’ve proven the fragile nature of road transportation systems regarding the recovery process of supply disruptions.
5 Fragility indicator and implications
Although road transportation systems have been mathematically proven to be intrinsically fragile, a quantitative approach to assess fragility across various networks is yet absent. It can also be of great interest to reveal factors contributing to or mitigating the fragile property of a road network. In Section 3, two generic methods, namely, second derivative and skewness for evaluating fragility have been discussed. Despite being proven to be positive and thus showcasing fragile, the exact value of the second derivatives varies even for different traffic states on the same MFD, let alone a cross-comparison among MFDs of various networks. Therefore, we resort to a skewness-based fragility indicator, so that one fixed value can be generated for a region based on the performance loss density distribution from multiple disruption recovery processes. For a given network, without considering hysteresis and other sources of uncertainties, a well-defined MFD can be produced and is widely utilized in the research of traffic control (Yildirimoglu and Geroliminis, 2014). Assuming no alteration in the driving behavior of the population, the well-defined MFD is primarily considered to be solely dependent on the network topology, including traffic control and signalization (Leclercq and Geroliminis, 2013). Therefore, in this section, we propose to develop a skewness-based indicator to evaluate the fragility of a given network, using only the MFD-related parameters obtained through analytical methods. This approach avoids the need for traffic sensor data, such as those from loop detectors, allowing for a pre-assessment of the network fragility early on during the design phase.
5.1 Empirical study on the skewness
As there is no direct way to define such a skewness-based indicator, we first design a unit MFD and empirically study the relationship between the skewness and the MFD-related parameters. Although typically more than three cuts can be generated to form an MFD through MoC, researchers sometimes simplify MFDs to be a trapezoidal shape with one forward, one stationary, and one backward cut (Tilg et al., 2020; Lee et al., 2023), which is sometimes also adopted for traffic control (Haddad and Geroliminis, 2012). In this section, we follow this simplification to facilitate creating a so-called unit MFD, similar to the isosceles FD proposed in Laval (2023), for the evaluation of fragility, formulated as $m=M(n,a_{f},a_{w},m_{\rm max},n_{\rm max})$, where $a_{f}$ and $a_{w}$ each denotes the forward cut and the backward cut, or $a_{N}$ and $a_{1}$ respectively. In addition, $m_{\rm max}$ represents the maximal trip completion derived from the stationary cut. As maximal vehicle accumulation, denoted as $n_{\rm max}$, is another indispensable parameter in MoC, for the proposed unit MFD, we fix $n_{\rm max}$ at $10000$ veh, resembling the $n_{\rm max}$ generated for the city center of Zurich in Fig. 12(b). An MFD for any given network can be scaled up or down to match the $n_{\rm max}$ of the unit MFD for cross-comparison. The real skewness of the network, denoted as $s$, is calculated as the asymmetry of the probability density distribution of TTS from multiple recovery processes with the same MFD but under different disruption magnitudes. With $\mu$ and $\sigma$ denoting the mean and standard deviation, the skewness can be formulated as:
$s=\frac{1}{N_{sample}}\sum_{i=1}^{N_{sample}}\left(\frac{TTS_{i}-\mu}{\sigma}\right)^{3}$ (47)
To ensure fair comparison, certain criteria have to be met. First, it is assumed that there is no base demand in the network. Second, when calculating TTS, the initial disruption vehicle accumulation $n^{\prime}$ follows a uniform distribution ranging between $5\%$ and $95\%$ of the maximal accumulation $n_{\rm max}$. The interval for vehicle increment for this setting is $50$, so a total number of $180$ data points are sampled. It is worth highlighting that to calculate skewness, a sufficient number of samples is required to reach satisfactory accuracy, and the chosen number of samples is substantially above the minimal threshold based on some pre-experiments. Also note these conditions merely serve as a reference instead of strict requirements, meaning that a range between $10\%$ and $90\%$ may also be used for determining the skewness. However, a larger range renders the skewness of different networks more distinguishable. Once the experiment setup is established, the aim is then to approximate the estimated skewness of a given network, denoted as $\tilde{s}$, based on its MFD-related parameters and a proposed approximation function $\tilde{s}=f(a_{f},a_{w},m_{\rm max})$.
One assumption for the analysis is $|a_{w}|\leq a_{f}$. As per Daganzo (1994), the backward wave travels several times slower than the free flow speed so the FD is formulated with a wave speed $|w|\leq u_{f}$. We also adopt the same assumption for the fragility indicator, although some further elaboration is needed. The FD is built on the analog with the hydrodynamic theory, where the well-known phenomenon of water hammer often occurs, which is a detrimental pressure surge when abruptly forcing to stop fluid motion in a rigid pipe (Ghidaoui et al., 2005). The pressure wave can propagate at a subsonic speed through the liquid along the pipe, which is considerably faster than the water flow speed, showing the possibility of backward wave speed being higher than the free flow speed. The major difference between traffic flow and water flow is the existence of safety distance in the former, but little gap in the latter. It can be easily proven that if the safety distance during driving could be the same as the effective vehicle length, the backward wave speed would equal the free flow speed. Therefore, it needs to be acknowledged that with the advancements in vehicular technologies such as Connected AVs (CAVs), which can form long platoons with minimal distance buffers, there is a likelihood that the backward wave speed could exceed the free flow speed in the future.
In Fig. 9, we present the heatmaps of $s$ as a function of $|a_{w}|$ and $a_{f}$ for different values of $m_{\rm max}$. Since the maximal vehicle accumulation $n_{\rm max}=10000$, which means a gradient $a_{f}\leq\cdot 10^{-4}$ will lead to the critical vehicle accumulation $n_{c}\geq n_{\rm max}$, and therefore, both axes representing $a_{f}$ and $|a_{w}|$ begin from $1.2\cdot 10^{-4}$ instead of $10^{-4}$ or even 00. The contour lines are also illustrated to indicate multiple levels of skewness. Several patterns can be observed based on the contour lines, which represent an implicit function between skewness and the MFD-related parameters $|a_{w}|=f(a_{f},m_{\rm max},s)$ instead of $s=f(a_{f},a_{w},m_{\rm max})$:
(a) $m_{\rm max}=0.5$
(b) $m_{\rm max}=0.75$
(c) $m_{\rm max}=1.0$
(d) $m_{\rm max}=1.25$
(e) $m_{\rm max}=1.5$
(f) $m_{\rm max}=1.75$
Figure 9: Skewness $s$ heatmap with $a_{f}$ and $a_{w}$ across different $m_{\rm max}$
Observation 1.
For the contour lines with the same skewness across different values of $m_{\rm max}$, it can be found that $m_{\rm max}$ has a scaling effect, which can be mathematically formulated as Eq. 48. For example, in Fig. 9(e) with $m_{\rm max}=1.5$, the contour line with $s=1.3$ crosses approximately the point $(a_{f},|a_{w}|)=(6.2,3.8)$, which means in Fig. 9(b) with $m_{\rm max}=0.75$, half of the $m_{\rm max}$ in Fig. 9(e), the same contour line $s=1.3$ will cross the point $(a_{f},|a_{w}|)=(3.1,1.9)$. This observation will be further discussed and validated later.
$\frac{|a_{w}|}{m_{\rm max}}=f(\frac{a_{f}}{m_{\rm max}},\tilde{s})$ (48)
Observation 2.
According to the assumption that $|a_{w}|\leq a_{f}$, we mainly focus on the upper triangular matrix. And for each skewness contour line within the upper triangular matrix, $|a_{w}|$ becomes a constant value when $a_{f}$ increases to infinity. However, it should be acknowledged that the contour lines with very low skewness in the lower triangular matrix do not specifically follow this observation. An exponential relationship can be roughly approximated between $|a_{w}|$ and skewness $\tilde{s}$ with coefficients $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$, as shown in Fig. 10(a), which can be mathematically formulated as:
$\lim_{a_{f}\to\infty}|a_{w}|=\beta_{1}e^{\beta_{2}(\tilde{s}-\beta_{3})}$ (49)
Observation 3.
For each skewness contour line, when $a_{f}$ gets closer to 00, the initial gradient of the contour line $\lim_{a_{f}\to 0}|\frac{a_{w}}{a_{f}}|$ will also become a constant. And when extended towards the upper left, they intercept exactly at the origin of the coordinates $(a_{f},|a_{w}|)=(0,0)$. The relationship between the gradient of the contour line $\lim_{a_{f}\to 0}|\frac{a_{w}}{a_{f}}|$ and skewness $\tilde{s}$ can be roughly approximated as a linear function, as shown in Fig. 10(b), which can be mathematically formulated as:
$\lim_{a_{f}\to 0}|\frac{a_{w}}{a_{f}}|=\beta_{4}\tilde{s}+\beta_{5}$ (50)
(a) Approximation of $\lim_{a_{f}\to\infty}|a_{w}|$
(b) Approximation of $\lim_{a_{f}\to 0}|\frac{a_{w}}{a_{f}}|$
Figure 10: Approximation of the coefficients
5.2 Skewness-based fragility indicator
The above observations suggest a significant relationship between the skewness and the MFD-related parameters worth further exploitation. Based on Observations 2 and Observation 3, these contour lines have been found to share great similarities with certain types of functions in the family of sigmoid curves. In this work, we call such approximators as activation functions $y=f_{\rm activation}(x)$, with the term borrowed from the deep learning domain, where the hyperbolic tangent $f_{\tanh}(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}$ is one of the most commonly applied functional forms. Likewise to Observation 2, the hyperbolic tangent function has a constant value at the right end $\lim_{x\to\infty}f_{\tanh}(x)=1$. And for Observation 3, a similar property can be observed that the gradient of the activation function at origin is also a constant $f^{\prime}_{\tanh}(0)=1$.
To scale the activation function to a certain skewness contour line according to these two observations to generate our proposed approximation function $f_{\rm approx}(\cdot)$, the below two conditions are to be satisfied:
$\lim_{a_{f}\to\infty}f_{\rm approx}(a_{f})=\lim_{a_{f}\to\infty}|a_{w}|\cdot\lim_{a_{f}\to\infty}f_{\rm tanh}(a_{f})$ (51a)
$\lim_{a_{f}\to 0}\frac{f_{\rm approx}(a_{f})}{a_{f}}=\lim_{a_{f}\to 0}|\frac{a_{w}}{a_{f}}|\cdot\lim_{a_{f}\to 0}\frac{f_{\rm tanh}(a_{f})}{a_{f}}$ (51b)
Therefore, the activation function for the approximation of skewness $s$ needs to be rewritten as:
$a_{w}$ $=f_{\rm approx}(a_{f})$
$=\lim_{a_{f}\to\infty}|a_{w}|\cdot f_{\rm tanh}\biggr{(}\frac{\lim_{a_{f}\to 0}|\frac{a_{w}}{a_{f}}|}{\lim_{a_{f}\to\infty}|a_{w}|}\cdot a_{f}\biggr{)}$ (52)
Now taking Observation 1 into account as well, and with coefficients $\beta_{1,...,5}$ being able to be approximated with Eq. 49, Eq. 50, and Fig. 10, the proposed approximation function is:
$\frac{a_{w}}{m_{\rm max}}=\beta_{1}e^{\beta_{2}(\tilde{s}-\beta_{3})}\cdot f_{\rm tanh}\biggr{(}\frac{\beta_{4}\tilde{s}+\beta_{5}}{\beta_{1}e^{\beta_{2}(\tilde{s}-\beta_{3})}}\cdot\frac{a_{f}}{m_{\rm max}}\biggr{)}$ (53)
Despite being an implicit function, the skewness $\tilde{s}$ can be solved with an iterative optimization algorithm, such as the Powell hybrid method or its modifications. Although with both ends $a_{f}\rightarrow 0$ and $a_{f}\rightarrow\infty$ of the contour lines being fixed, the hyperbolic tangent function alone does not necessarily guarantee a rather decent approximation. Hence, we introduce some other activation functions that resemble the hyperbolic tangent function with a list below, which will be applied and cross-compared for the approximation of the skewness of a network. All these activation functions share the same characteristics that $\lim_{x\to\infty}f_{\rm activation}(x)=1$ and $f^{\prime}_{\rm activation}(0)=1$, but with varying curvature to reach $a_{f}\rightarrow\infty$:
  • 1.
[…]
  • 2.
    Gudermannian function: $f_{\rm gd}(x)=2\arctan(\tanh(\frac{x}{2}))$
  • 3.
    Arctangent function: $f_{\rm arctan}(x)=\frac{2}{\pi}\arctan(\frac{2}{\pi}x)$
  • 4.
    Inverted square root unit: $f_{\rm isru}(x)=\frac{x}{\sqrt{1+x^{2}}}$
  • 5.
    A generalized form of the inverted square root unit: $f_{\kappa}(x)=\frac{x}{(1+|x|^{\kappa})^{1/\kappa}}$
5.3 Validation
To validate the effectiveness of our proposed skewness-based fragility indicator, the approximation error $\Delta s$ between the real skewness $s$ from the unit MFD and the approximated skewness $\tilde{s}$ of each activation function is computed and presented in Fig. 11 with $m_{\rm max}=1$. In Fig. 11(a) we select $k=5$ to represent the generalized form of the inverted square root unit. A darker shape of red or blue indicates the skewness from the approximation $\tilde{s}$ is greater or smaller than the real value $s$. It can be observed that the activation functions $f_{\kappa=5}$ and $f_{\rm erf}$ show seemingly superior performance, with the vast majority of the heatmap cells lying below a $|\Delta s|$ of $0.1$. It should also be noted that at the upper margin of the heatmap with $a_{f}\gg|a_{w}|$, particularly for $f_{\kappa=5}$ and $f_{\rm erf}$, a red strip with significant positive error can be observed. We acknowledge but accept the rather large error in the red strip, for the same reason as the lower triangular matrix $a_{f}<|a_{w}|$, this strip with $a_{f}\gg|a_{w}|$ lacks realisticity either.
(a) $f_{\kappa=5}$
(b) $f_{\rm erf}$
(c) $f_{\rm tanh}$
(d) $f_{\rm gd}$
(e) $f_{\rm isru}$
(f) $f_{\rm arctan}$
Figure 11: Error of approximation by applying different activation functions with $m_{\rm max}=1$
To better quantify the error $\Delta s$ from different approximation functions, we summarize the error into Table 3 with Mean Absolute Error (MAE), Mean Squared Error (MSE), and Root Mean Square Error (RMSE). Other values of $\kappa$ for the generalized form of the inverted square root unit $f_{\kappa}$ are also presented. The activation function $f_{\kappa=5}$ showcased the best performance, with an MAE and an RMSE being $0.032$ and $0.043$ respectively. Considering the skewness in the upper triangular matrix in Fig. 9(c) varies roughly between $0.7-1.5$ and shows an average skewness of $1.136$, the MAE and RMSE demonstrate an error well below 5% and thus validate the accuracy of our proposed approximation function. With $f_{\kappa=5}$, we also showcase the approximation accuracy for varying values of $m_{\rm max}$ in Table 3. While the error still lies below $5\%$ with $m_{\rm max}$ larger than $1$, the accuracy deteriorates when $m_{\rm max}$ becomes smaller. This is partially attributable to the function form of approximation in Eq. 49 and Eq. 50 as well as the coefficients $\beta_{1,...,5}$. The current coefficients are approximated with the data points from a trimmed range in Fig. 10 to achieve the least error in the case of $m_{\rm max}=1$. If a slightly larger range is considered, a rather uniform value of RMSE at around $0.75-0.85$ can be observed regardless of the different values of $m_{\rm max}$, roughly $7\%$ of the average skewness, validating the Observation 1. Also, for the approximation of $\lim_{a_{f}\to\infty}|a_{w}|$, a polynomial can be applied in replacement of the current exponential function to achieve potential better performance.
Table 3: Error of approximation of applying different activation functions
Function MAE MSE RMSE With $f_{\kappa=5}$ MAE MSE RMSE
$f_{\rm erf}$ 0.037 0.0023 0.048 $m_{\rm max}=0.5$ 0.120 0.034 0.185
$f_{\rm tanh}$ 0.044 0.0033 0.057 $m_{\rm max}=0.75$ 0.044 0.0045 0.067
$f_{\rm gd}$ 0.054 0.0045 0.067 $m_{\rm max}=1.0$ 0.032 0.0018 0.043
$f_{\rm isru}$ 0.084 0.0089 0.094 $m_{\rm max}=1.25$ 0.034 0.0023 0.048
$f_{\rm arctan}$ 0.152 0.0261 0.162 $m_{\rm max}=1.5$ 0.038 0.0026 0.051
$f_{\kappa=4}$ 0.033 0.0019 0.043 $m_{\rm max}=1.75$ 0.040 0.0028 0.053
$f_{\kappa=5}$ 0.032 0.0018 0.043
$f_{\kappa=6}$ 0.032 0.0019 0.044
5.4 Implications
Aside from the proposed approximation function to estimate the skewness-based fragility indicator, insights into the fragility together with the MFD-related parameters can also be drawn from the heatmaps in Fig. 9. When $|a_{w}|$ is not sufficiently large, an increase in $a_{f}$ contributes to the network’s fragility, whereas a larger $|a_{w}|$ reduces such fragility. The implications are twofold:
    To minimize urban traffic accident rates and reduce noise pollution, aligned with the long-term goal of Vision Zero in the EU and worldwide to reach no traffic fatalities and severe injuries (Doecke et al., 2020), many cities have implemented stricter speed limits, such as the Tempo 30 regulation in Zurich (Menendez and Ambühl, 2022). Although this may reduce the overall serviceability of the urban road networks, it enhances the city’s antifragility when dealing with traffic disruptions.
    While backward wave speed is generally linked to the driving behavior of human-driven vehicles, Makridis et al. (2024) recently found that adaptive cruise control, a precursor to AVs, can increase backward wave speed with minimum headway settings (although still not as fast as free-flow speed), suggesting that introducing AVs into transportation systems could potentially improve the network’s antifragility. However, this enhancement of antifragility at the macroscopic level may result in fragility at the microscopic level, as minimum safety distances can cause propagating string instability under perturbations (Makridis et al., 2021), if the connectivity and cooperation levels are not sufficiently high.
@@ 6 Numerical simulation
Even though researchers generally consider MFDs to be well-defined, road transportation systems in the real world and their MFDs are always subject to stochasticity all the time, as shown in Geroliminis and Daganzo (2008); Ambühl et al. (2021), and this is the same case for FDs as well (Siqueira et al., 2016; Qu et al., 2017). Therefore, when validating a newly proposed traffic control algorithm, it has become a common practice to account for model uncertainties and showcase the method’s robustness, such as in Geroliminis et al. (2013); Haddad and Mirkin (2017); Zhou and Gayah (2023). In our study, however, the model stochasticity cannot be directly reflected in the mathematical analysis, hence, it is indispensable to show the influence of realistic stochasticity on the fragile nature of transportation systems with a numerical simulation, i.e., whether the system still maintains the same fragile response under real-world errors in MFD when a demand or supply disruption is present.
6.1 Experimental setup with real-world parameters
@@
In this section, we simulate the disruption recovery process. The MFD of the studied region is generated by applying MoC following Daganzo and Geroliminis (2008) with realistic parameters in the city center of Zurich. Some parameters, e.g., free-flow speed, back-propagation speed, maximal density, and capacity are provided in Ambühl et al. (2020) for Zurich with queried routes in Google API and with other validation methods. The total and average lane length for the city center is determined through SUMO with OpenStreetMap API. The average trip length of Zurich is studied in Schüssler and Axhausen (2008). We introduce stochasticity in the city center of Zurich with real traffic light data, which is publicized by the Statistical Office of Zurich and accessible in Genser et al. (2023, 2024). The authors acknowledge that MoC is developed with the premise of a homogeneous region, and given that the available data is limited to only one main intersection in this region, we assume that this intersection serves as a representative sample for the city center region. Also, since the signalization in Zurich is actuated based on the present traffic flow, they do not strictly follow a fixed-time signal cycle. Despite this actuation, a concentrated distribution can be easily observed in the dataset and we assume the green split of the cycle follows a normal distribution. The offset is presumed zero as the assumption for the actuated signal in Yokohama in Daganzo and Geroliminis (2008). According to the daily average traffic density of Zurich in Ambühl et al. (2021), we approximate the traffic demand, which is also the trip completion when the traffic state is at equilibrium, is about $0.6$ veh/s for our studied region. This corresponds to an accumulation of around 975 vehicles in the city center. The parameters and the related values are summarized in Table 4.
Table 4: Real-world parameters for the city center of Zurich
Parameters Notation Unit Value
Free-flow speed $u_{l}$ $\rm m/s$ 12.5
Backward wave speed $w_{l}$ $\rm m/s$ 6.0
Maximal density $k_{\rm max}$ $\rm veh/m$ 0.145
Capacity $c_{l}$ $\rm veh/s$ 0.51
Total lane length $D$ $\rm m$ 68631
Average lane length $l$ $\rm m$ 167
Average trip length $L$ $\rm m$ 7110
Signal cycle time $C$ $\rm s$ 50
Signal green time (mean) $\mu_{G}$ $\rm s$ 14.8
Signal green time (std.) $\sigma_{G}$ $\rm s$ 2.5
Offset $\delta$ $\rm s$ 0
Traffic demand $m_{0}$ $\rm veh/s$ 0.6
The average green time of the signal is 14.8 s and its standard deviation is 2.5 s. With an interval of one standard deviation, i.e., a set of green time $\mathcal{G}=\{\mu_{G}-\sigma_{G},\mu_{G},\mu_{G}+\sigma_{G}\}$, and following the MoC described in Daganzo and Geroliminis (2008), we can produce stationary, forward, and backward cuts in gray, blue, and red lines, for each of the MFD with a unique green time from $\mathcal{G}$, as Fig. 12(a) shows with dotted lines, solid lines, and dashed lines. The MFD with a longer green time $\mu_{G}+\sigma_{G}$ yields a greater MFD and vice versa. The MFDs share the same maximal vehicle density of around 0.145 veh/m, which corresponds to a gridlock accumulation of about 10000 vehicles for the studied region based on Eq. 2a and Eq. 2b. In the following simulations for demand and supply disruptions with realistic model stochasticity, we study the congestion dissipation processes in the road network as traffic recovers from disruptions. At each timestep, a stochastic signal green time is sampled following normal distribution $G\sim\mathcal{N}(14.8,2.5)$ based on real-world data, leading to an uncertain MFD profile as Fig. 12(b) shows. The green scattering points connected with thin dotted lines display an example of a congestion dissipation process with an initial disruption of about 7000 vehicles and are sampled from every $100^{th}$ point for clarity.
(a) The MFD of the city center of Zurich through MoC
(b) Stochasticity on the MFD
Figure 12: The MFD of the city center of Zurich through MoC and stochasticity
6.2 Demand disruptions